OpenStudy (anonymous):

Please help... Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at y = ± 5/6x

OpenStudy (mertsj):

I see the center is (0,0). Do you see that?

OpenStudy (anonymous):

yes but how does that fit in the equation?

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1 \]

OpenStudy (mertsj):

Center is (h,k). That is the equation of a hyperbola with horizontal transverse axis. Is the transverse axis of your hyperbols horizontal?